OpenStudy (anonymous):

write an equation of an ellipse in standard form with the center at the origin and with the given characteristics. vertex at (-5,0) and co-vertex at (0,4)

OpenStudy (anonymous):

hey Einstein i thought you died in 1955

OpenStudy (anonymous):

Rebirth!!

OpenStudy (anonymous):

but I need help

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

@kiwi_is_here_to_help can you help me?

OpenStudy (anonymous):

@Aiko can you help me

OpenStudy (anonymous):

@ali1029

OpenStudy (anonymous):

i think it would be x^2/25+y^2/16=1

OpenStudy (anonymous):

do you have time to explain? @ali1029

OpenStudy (anonymous):

sure :)

OpenStudy (jdoe0001):

the vertex of an ellipse, will be where the MAJOR AXIS is at the covertex, will be where the MINOR AXIS is at we know that the center of this ellipse is at the origin so, what do you think is the "a" component and "b" components?

OpenStudy (anonymous):

-5 and 4 @jdoe0001

OpenStudy (jdoe0001):

 what about now?

OpenStudy (anonymous):

5 and 4 will be our a^2 b^2. when y=0, x=5, x^2/25 + y^2/16 = 1. put the 5 under the x parts, and the 4 under the y parts, since our options for a and b are 5 and 4 ..... and when y=0, x =-5 .... therefore we need a=5 and b=4, the standard setup squares those values into a^2 = 25 and b^2 = 16

OpenStudy (jdoe0001):

the "a" component, is the distance from the center, to the vertex the "b" component, is the distance from the center, to the covertex

OpenStudy (anonymous):

I'm somewhat understanding it though, @jdoe0001 whats your explanation